OpenStudy (anonymous):

I've tried to answer this, but I just don't seem to get the right equation to begin with. Help, please? The distance between city A and city B is 280 km. If in going from city A to B, a car traveled 30 kph faster than usual, the trip would've required 3 hours less than the usual time. Find the original speed of the car. Thank you!! <3

OpenStudy (anonymous):

\[\frac{280}{u}-\frac{280}{u+30}=3\]

OpenStudy (anonymous):

Hi! Could you also explain how you came up with the equation? Thank you! :)

OpenStudy (anonymous):

distance/speed=time

OpenStudy (anonymous):

original speed=\[40 km/h\]

OpenStudy (anonymous):

since distance = speed * t so... Time=d/s so,(280/x) - (280/x+30)= 3 so, Time - Time = Time :)

OpenStudy (anonymous):

Hi phoebeco. understood ?? Medel plz :) :)

OpenStudy (anonymous):

distance/original speed=time taken before increasing speed. distance/increased speed=time taken after increasing speed subtraction yields 3hrs. as per question.

OpenStudy (anonymous):

plz contribute to my medel.... have none :( inspire me to help by giving medel

OpenStudy (anonymous):

hope u r inspired:)

OpenStudy (anonymous):

:) yes...........

OpenStudy (anonymous):

thank you guys, so much!! :) well explained!